OpenStudy (anonymous):

Write an equation for the following problems and solve. The sum of a number and 2 is 6 less than twice that number.

OpenStudy (anonymous):

Use a varible for the mysterious number!!!

OpenStudy (anonymous):

6(2x)-x+2=? i think

OpenStudy (anonymous):

it would be (x+2) - 6 = 2x

OpenStudy (anonymous):

x + 2 is a sum of a number and 2

OpenStudy (anonymous):

2x is doulbe that number

OpenStudy (anonymous):

and - 6 makes it 6 less!

OpenStudy (anonymous):

funfufunfunfun

OpenStudy (anonymous):

let the number be x x+2= 6-2x 3x=4 x=4/3

OpenStudy (anonymous):

which one is it????

OpenStudy (anonymous):

\[x+2=6\] and \[6<2x\] and dividing through by 2, \[3<x\] therefor \[x+2=6 \left\{ x: x>3 \right\}\]

OpenStudy (anonymous):

"The sum of a number" x "and 2" x+2 "is" x+2 = "6 less" x+2 = -6 "than twice the number" x+2 = -6 + 2x x = 8